OpenStudy (anonymous):

Multiply out and simplify to obtain a sum of product: (A+B)(C+B)(D'+B)(ACD'+E)

OpenStudy (anonymous):

Boolean algebra

OpenStudy (neer2890):

(A+B)(C+B)(D'+B)(ACD'+E) =(AC+AB+BC+B)(D'+B)(ACD'+E) =(AC+B)(D'+B)(ACD'+E) [AB+BC+B=AB+B(C+1)=AB+B=B(A+1)=B] =(ACD'+ACB+BD'+B)(ACD'+E) =(ACD'+B)(ACD'+E) [ACB+BD'+B=ACB+B(D'+1)=ACB+B=B(AC+1)=B] =ACD'+ACD'E+ACD'B+BE =ACD'+BE